# Supervising 3D Talking Head Avatars with Analysis-by-Audio-Synthesis

### v1

@@ 5.1 Datasets
The selection of datasets for THUNDER is challenging. On the one hand, 4D scan datasets such as VOCASET [12] provide good reconstructions but are too small to train generative models. On the other hand, in order to have a high-enough quality geometry from which speech can be predicted, we cannot rely on in-the-wild datasets, as reconstruction artifacts decrease the quality of the GT and hence our mesh-to-speech and talking head models.
Hence, we use a collection of public in-the-lab captured video datasets commonly used for lip-reading, namely TCD-TIMIT [24], RAVDESS [42] and GRID [10], totaling 117 subjects, uttering 42140 short sentences, totaling 3535816 frames at 25FPS (over 39 hours). These datasets provide clear imagery without in-the-wild complexity and can be reconstructed by off-the-shelf methods to a satisfactory degree. Following EMOTE [14], we utilize the publicly available INFERNO framework [15], which contains a SOTA face reconstruction system [80, 13, 21]. We split each of the datasets by subjects, 70% for training and 15% each for validation and testing.
@@ 5.2 Mesh-To-Speech
Implementation details. We initialize the conformer and prediction heads with pretrained weights from Choi et al. [8]. The input MLP is trained from scratch. We train the model with the Adam optimizer [38] for 20 epochs and select the checkpoint with the lowest validation loss (reached within 5 epochs). The training batch size is 16 and the input sequence is trimmed to 125 frames.
Quantitative comparison. Since the mesh-to-speech model is the first of its kind, we compare mesh-to-speech to a recent silent-video-to-speech model [8] which we finetune on the videos from our training dataset for fair comparison. Note that FLAME does not model the teeth or tongue, which makes the mesh-to-speech task arguably harder than video-to-speech. Teeth and tongue, which may be visible on video, are critical for production of certain phonemes, such as labiodentals (/f/, /v/), or alveolar phonemes (/s/, /z/, /t/, /d/, /l/) [39]. Consequently, we do not expect mesh-to-speech to match the performance of a video-based method.
Table 1 reports the standard video-to-speech metrics, namely Short-Time Objective Intelligibility [61], Extended STOI [30] to measure the intelligibility of the generated samples, Perceptual Evaluation of Speech Quality (PESQ) [57] (narrow and wide band) to measure the perceptual quality, and Word Error Rate (WER), which we measure with the SOTA Whisper ASR network [54]. As expected mesh-to-speech is not as accurate as video-to-speech, but the performance is comparable, suggesting that there is sufficient audio signal in the 3D to be useful for our downstream task. Note that mesh-to-speech is, however, more accurate than Choi et al.’s original (but not the fine-tuned) model. We experiment with the following input spaces for mesh-to-speech - selection of mouth vertices $\mathbf{V}_{m\in\mathcal{M}}$ (mouth2s), all vertices $\mathbf{V}$ (exp2s), and FLAME expression parameters $\boldsymbol{x}$ (exp2s) and report the results in Tab. 1. Remarkably, the model which takes global expression vectors, performs slightly better than the selection of mouth vertices $\mathbf{V}_{m\in\mathcal{M}}$. Despite that, we choose mouth2s to be our final model, as it results in better lip-sync supervision (discussed in Sec. 5.3 and Tab. 3). Please refer to the Sup. Video for audible comparison and Sup. Mat. for further discussion on performance.
@@
Table 1: Quantitative comparison of video-to-speech and various forms of mesh-to-speech (mouth2speech, face2speech and exp2speech).
@@ 5.3 Speech-Driven Facial Animation
Baselines. We compare THUNDER to other speech-driven avatar methods that are also trained on pGT and which output 3DMM parameters (as opposed to vertices). We reimplementated several SOTA methods, namely FlameFormer (a re-implementation of FaceFormer [19]), Media2Face [78] and DiffPoseTalk [60] (see Sup. Mat. for details). These reimplementations are trained on our data and are labeled with an asterisk*. Unless stated otherwise, we keep the pretrained Wav2Vec weights frozen. Models that finetune Wav2Vec are denoted with a suffix -T (T for trainable).
EMOTE 1.06 0.221 0.442 0.466 0.0563 0.0401 0.248 0.24 0.0141 0.104
THUNDER 0.802 0.426 0.639 0.29 0.0322 0.04 0.134 0.241 0.0122 0.0806
Table 2: Quantitative comparison of THUNDER with other speech-driven avatar methods. Asterisk* indicates we have re-implemented the baseline and trained it on our dataset. (1) On lip-sync. THUNDER outperforms all baselines on lip-sync metrics (LVE, CCC, PCC, DTW). (2) On expression diversity. DiffPoseTalk exhibits higher temporal face diversity T-DIV-U and L-DIV-U, likely due to the training data reconstructed from videos in-the-wild, which contains more dynamic motions than our in-the-lab reconstructions. EMOTE exhibits high sample upper face diversity S-DIV-U and expression diversity S-DIV-EXP which is due to the random sampling of speaking/emotion/intensity styles, which while high, may not be fitting to the particular input audio. THUNDER w/o m2s and THUNDER follow next in S-DIV-U.
Evaluation protocol. For each test audio sequence, we generate 32 outputs. Each of these is initialized with a different random noise in the case of diffusion models, or randomly sampled subject-id condition in case of EMOTE and FlameFormer. Then we compute the following evaluation metrics.
Lip Vertex Error. Following [56, 60, 71] we report LVE, which calculates the maximum L2 error of all lip vertices for each frame and then computes the average over all frames.
Dynamic Time Warping. We also report the DTW proposed by Thambiraja et al. [66]. First, distances between the mid-points of lower and upper lips are calculated for both the predictions and GT. The two resulting time series are then used to compute the DTW distance.
@@
Face dynamic deviation. An important aspect of facial animation, is the expressiveness of the upper face vertices. Following CodeTalker [71], we report Face Dynamics Deviation (FDD), which measures the differences between the temporal standard deviation of the pGT and the predicted vertices, averaged over a set of vertices. We report upper-FDD (FDD-U) and lip-FDD (FDD-L).
Sample diversity. One of the most important aspects of stochastic models is that they generate diverse animations for the same input. Remarkably, previous work has not analyzed this [59, 60, 78]. To measure this, we compute standard deviations of vertex distances between the pGT and the predictions for the same input at frame $t$. Formally, given a tensor of per-vertex L2 distances $\mathbf{D}\in\mathbb{R}^{N\times S\ \times T\times n_{v}}$, with $N$ being the size of the test dataset and $S=32$ is the number of sampled outputs per input. We compute the standard deviation along the sample dimension $S$. Other dimensions are subsequently averaged. We measure sample diversity for the upper face region (S-Div-U) and the lip region (S-Div-L).
@@ 5.3.1 Quantitative evaluation
Comparison to other methods. In Tab. 3 we compare recent methods, trained on pseudo-GT, namely EMOTE [14] (deterministic SOTA) and DiffPoseTalk [60] (stochastic SOTA). Additionally, we include a FlameFormer trained on our data, and THUNDER trained without mesh-to-speech. Please note that for fair comparison, the original DiffPoseTalk is run without the style condition, as to not limit its diversity.
Mesh-to-speech loss with other methods. Tab. 2 shows that mesh-to-speech is applicable to other methods. We re-implement three methods that incorporate additional conditioning mechanisms and train them with and without our mesh-to-speech loss. Our new loss improves performance on the lip-sync metrics. For visual results, refer to Sup. Video.
Experiment Name LVE $\downarrow$ L-CCC $\uparrow$ L-PCC $\uparrow$ DTW $\downarrow$ S-DIV-U $\uparrow$ T-DIV-U $\uparrow$ S-DIV-L $\downarrow$ T-DIV-L $\uparrow$ T-FDD-U $\downarrow$ T -FDD-L $\downarrow$
Mesh-to-speech with other methods FlameFormer* 0.809 0.368 0.57 0.291 0.0271 0.0372 0.132 0.239 0.0117 0.0909
@@
Mesh-to-speech input space ablation THUNDER w/o m2s 0.879 0.359 0.568 0.329 0.0419 0.044 0.21 0.254 0.0118 0.0932
@@
Table 3: Mesh-To-Speech applied to other methods. The application of mesh-to-speech to our reimplementations of Media2Face (diffusion with image conditioning), DiffPoseTalk (diffusion with style feature conditioning) and FlameFormer (deterministic with one-hot conditioning) improves lip-sync metrics at slight expense of expression diversity. Ablation of the mesh-to-speech model. All of the mesh-to-speech variants are applicable. Mouth2s results in the best lip-sync performance and exp2s scores best on diversity metrics, while face2s scores the worst on diversity.
@@
Ablation study. We ablate our design decisions. (1) The input space for the mesh-to-speech loss (see Tab. 3), (2) the weight of the mesh-to-speech loss, and (3) the trainability of the input audio encoder (see Sup. Mat. and video).
@@
Disentanglement effect. To test whether mesh-to-speech helps preserve the effect of good lip-sync in the presence of other editing conditions, we additionally run our reimplementations of Media2Face with conditioning inputs extracted from emotional images. For the results, see Sup. Mat.
@@
Perceptual study. We conduct a perceptual study in which we compare THUNDER with other methods. This includes original implementations of DiffPoseTalk and EMOTE, our implementation of FlameFormer, and THUNDER without mesh-to-speech. The participants are shown two videos side-by-side and asked to rate animations on a five-point Likert scale (strong/weak preference for A or B or indifferent). We ask the participants to rate three aspects of the animation: lip-sync quality, dynamism, and realism. THUNDER outperforms nearly all other methods on lip-sync, including those with the trainable audio encoder (denoted -T) with the exception of the original DiffPoseTalk implementation (see Fig. 4 and Fig. 8 in Sup. Mat.). We attribute the performance of DiffPoseTalk to its dataset (TFHP), which contains unscripted and more natural animations. To verify this, we also compare DiffPoseTalk with the GT (last row). Remarkably, the participants preferred DiffPoseTalk over GT. For the remaining results of the study and further discussion, see Sup. Mat.
@@
Figure 4: Perceptual study results of THUNDER (ours) compared with other methods. We compare THUNDER-F with methods having both a trainable audio encoder (DiffPoseTalk, EMOTE, THUNDER-T w/o m2s) and frozen encoders (FlameFormer-F, THUNDER-F w/o m2s).
@@ 5.4 Limitations
THUNDER has the following limitations. The absence of teeth and tongue inherently creates ambiguity, making it more difficult to produce high quality audio compared to silent video-to-speech. Further, our pseudo-GT is not a perfect reconstruction but has occasional artifacts or inaccuracies. Also, our current mesh-to-speech architecture is fully deterministic and hence is not completely capable of capturing the prediction ambiguities. All of these may result in suboptimal effects of the audio cycle-consistency loss. Employing a large-scale dataset of high quality 3D scans, a face model that models teeth and tongue, and upgrade mesh-to-speech to a stochastic model may boost the benefits of the mesh-to-speech loss. Finally, the inference process of the diffusion model is currently computationally intensive. Future work should address this by employing more efficient solvers such as DPM++ [43].
@@ 6 Conclusion
Appendix A Mesh-to-Speech
A.1 Per-Dataset Performance Analysis.
The performance of both mesh-to-speech and video-to-speech models varies dramatically depending on the dataset. Datasets with limited number of utterances and vocabulary such as RAVDESS or GRID are less challenging, and the models can regress relatively accurately, sometimes to the point of perfect intelligibility. On the more challenging TCD-TIMIT dataset, which contains more complex vocabulary, words are often missed. However, despite the relatively high word error rate (WER) on TCD-TIMIT, the produced sounds, remain very realistic, given the animation. Please see Tab. 4 for per-dataset breakdown and refer to the supplementary video for an audible comparison.
@@ Appendix A Mesh-to-Speech
Appendix B Speech-Driven Animation
B.1 Architecture of Baselines.
Media2Face. Our reimplementation of Media2Face is built on our diffusion architecture. In training, we provide additional input of ImageBind[22] features extracted from images of the corresponding videos of the training set. Like the audio condition, this condition is dropped 20% of the time.
DiffPoseTalk. Our reimplementation of DiffPoseTalk is also built on our diffusion architecture. It is trained in two stages. In the first stage we train the contrastive style encoder as proposed by the authors [60]. In training of DiffPoseTalk, we provide the additional style feature on the input. The style feature is extracted from the GT animation sequence. Like the audio condition, the style condition is dropped 20% of the time.
Figure 6: Architecture of re-implemented baselines. (a) Our version of Media2Face [78] utilized the same architecture as THUNDER but takes an extra ImabeBind feature on the input. (c) Our re-implementation of DiffPoseTalk [60] is also based on THUNDER and takes an extra style feature from a pretrained style encoder. (c) FlameFormer is an adapted FaceFormer [19]. FlameFormer predicts FLAME expression instead of full vertex space. Furthermore, FlameFormer adapts a non-autoregressive BERT-like prediction mechanism (akin to EMOTE [14]).
FlameFormer. Our FlameFormer architecture is based on FaceFormer. Like FaceFormer, FlameFormer is a transformer-based deterministic speech-driven animation network that takes addition categorical style condition on the input. Instead of predicting the full vertex space like the original FaceFormer does, FlameFormer predicts FLAME 3DMM coefficients $\boldsymbol{x}$. Additionally, instead of of using the autoregressive loop with a transformer decoder, FlameFormer’s decoder is a transformer encoder, which eliminates the need for the autoregressive formulation. This does not degrade performance and makes the model more efficient, as was already shown by Daněček et al. [14], which have proposed the FlameFormer baseline in their paper. Furthermore, we extend FlameFormer’s conditioning inputs - the subject identity (one-hot vector) is complemented with a one-hot vector for 8 basic emotions and three intensities.
B.2 Experiments
B.2.1 Disentanglement

### v2

@@ 5.1 Datasets
The selection of datasets for THUNDER is challenging. On the one hand, 4D scan datasets such as VOCASET [cudeiro2019capture] provide good reconstructions but are too small to train generative models. On the other hand, in order to have a high-enough quality geometry from which speech can be predicted, we cannot rely on in-the-wild datasets, as reconstruction artifacts decrease the quality of the GT and hence our mesh-to-speech and talking head models.
Hence, we use a collection of public in-the-lab captured video datasets commonly used for lip-reading, namely TCD-TIMIT [Harte2015tcdtimit], RAVDESS [livingstone_2018_1188976] and GRID [cooke2016grid], totaling 117 subjects, uttering 42140 short sentences, totaling 3535816 frames at 25FPS (over 39 hours). These datasets provide clear imagery without in-the-wild complexity and can be reconstructed by off-the-shelf methods to a satisfactory degree. Following EMOTE [danecek2023emote], we utilize the publicly available INFERNO framework [inferno2023], which contains a SOTA face reconstruction system [MICA:ECCV2022, EMOCA:CVPR:2021, filntisis2022visual]. We split each of the datasets by subjects, 70% for training and 15% each for validation and testing.
@@ 5.2 Mesh-To-Speech
Implementation details. We initialize the conformer and prediction heads with pretrained weights from Choi et al. [choi2023lip2speechunit]. The input MLP is trained from scratch. We train the model with the Adam optimizer [Kingma2015] for 20 epochs and select the checkpoint with the lowest validation loss (reached within 5 epochs). The training batch size is 16 and the input sequence is trimmed to 125 frames.
Quantitative comparison. Since the mesh-to-speech model is the first of its kind, we compare mesh-to-speech to a recent silent-video-to-speech model [choi2023lip2speechunit] which we finetune on the videos from our training dataset for fair comparison. Note that FLAME does not model the teeth or tongue, which makes the mesh-to-speech task arguably harder than video-to-speech. Teeth and tongue, which may be visible on video, are critical for production of certain phonemes, such as labiodentals (/f/, /v/), or alveolar phonemes (/s/, /z/, /t/, /d/, /l/) [Ladefoged2011phonetics]. Consequently, we do not expect mesh-to-speech to match the performance of a video-based method.
Table 1 reports the standard video-to-speech metrics, namely Short-Time Objective Intelligibility [Taal2011stoi], Extended STOI [Jensen2016estoi] to measure the intelligibility of the generated samples, Perceptual Evaluation of Speech Quality (PESQ) [Rix2001pesq] (narrow and wide band) to measure the perceptual quality, and Word Error Rate (WER), which we measure with the SOTA ASR network Whisper[radford2023whisper]. As expected, mesh-to-speech is not as accurate as video-to-speech, but the performance is comparable, suggesting that there is sufficient audio signal in the 3D to be useful for our downstream task. Note that mesh-to-speech is, however, more accurate than Choi et al.’s original (but not the fine-tuned) model. We experiment with the following input spaces for mesh-to-speech - selection of mouth vertices $\mathbf{V}_{m\in\mathcal{M}}$ (mouth2s), all vertices $\mathbf{V}$ (exp2s), and FLAME expression parameters $\boldsymbol{x}$ (exp2s) and report the results in Tab. 1. Remarkably, the model which takes global expression vectors, performs slightly better than the selection of mouth vertices $\mathbf{V}_{m\in\mathcal{M}}$. Despite that, we choose mouth2s to be our final model, as it results in better lip-sync supervision (discussed in Sec. 5.3 and Tab. 2). Please refer to the Sup. Video for audible comparison and Sec. 7.1 in Sup. Mat. for further discussion on performance.
@@
Table 1: Quantitative comparison of video-to-speech and forms of mesh-to-speech (mouth2speech, face2speech and exp2speech).
@@ 5.3 Speech-Driven Facial Animation
Baselines. We compare THUNDER to other speech-driven avatar methods that are also trained on pGT and which output 3DMM parameters (as opposed to vertices). We reimplemented and trained several SOTA methods, namely FlameFormer (an adaptation of FaceFormer [fan2022faceformer]), Media2Face [zhao2024media2face] and DiffPoseTalk [sun2024diffposetalk] (see Sup. Mat. for details). Unless stated otherwise, we keep the pretrained Wav2Vec weights frozen. Models that finetune Wav2Vec are denoted with a suffix -T (T for trainable).
Evaluation protocol. For each test audio sequence, we generate S=32 outputs. Each of these is initialized with a different random noise in the case of diffusion models, or randomly sampled subject-id condition in case of FlameFormer. Then we compute the following evaluation metrics, averaged out across the S outputs (unless stated otherwise).
Lip Vertex Error. Following [richard2021meshtalk, sun2024diffposetalk, xing2023codetalker] we report LVE, which calculates the maximum L2 error of all lip vertices for each frame and then computes the average over all frames.
Dynamic Time Warping. We also report the DTW proposed by Thambiraja et al. [bala2022imitator]. First, distances between the mid-points of lower and upper lips are calculated for both the predictions and GT. The two resulting time series are then used to compute the DTW distance.
@@
Face dynamic deviation. An important aspect of facial animation, is the expressiveness of the upper face vertices. Following CodeTalker [xing2023codetalker], we report Face Dynamics Deviation (FDD), which measures the differences between the temporal standard deviation of the pGT and the predicted vertices, averaged over a set of vertices. We report upper-FDD (FDD-U) and lip-FDD (FDD-L).
Sample diversity. One of the most important aspects of stochastic models is that they generate diverse animations for the same input. Remarkably, previous work has not analyzed this aspect [stan2023facediffuser, sun2024diffposetalk, zhao2024media2face]. We compute standard deviations of vertex distances between the pGT and the predictions for the same input at frame $t$. Formally, given a tensor of per-vertex L2 distances $\mathbf{D}\in\mathbb{R}^{N\times S\ \times T\times n_{v}}$, with $N$ being the size of the test dataset and $S=32$ is the number of sampled outputs per input. We compute the standard deviation along the sample dimension $S$. Other dimensions are subsequently averaged. We measure sample diversity for the upper face region (S-Div-U) and the lip region (S-Div-L).
@@ 5.3.1 Quantitative evaluation
Tab. 2 reports the computed metrics computed on the test set for THUNDER, baselines and ablations.
Experiment Model LVE $\downarrow$ L-CCC $\uparrow$ L-PCC $\uparrow$ DTW $\downarrow$ S-DIV-U $\uparrow$ T-DIV-U $\uparrow$ S-DIV-L $\downarrow$ T-DIV-L $\uparrow$ T-FDD-U $\downarrow$ T -FDD-L $\downarrow$
THUNDER and mesh-to-speech input space THUNDER w/o m2s 0.879 0.359 0.568 0.329 0.0419 0.044 0.21 0.254 0.0118 0.0932
THUNDER and trainable audio enc. THUNDER-T w/o m2s 0.723 0.428 0.623 0.266 0.02 0.0411 0.0656 0.216 0.0118 0.0811
THUNDER-T w/ mouth2s 0.709 0.445 0.66 0.256 0.021 0.039 0.0669 0.202 0.0118 0.0788
Mesh-to-speech with other methods and additional input conditions FlameFormer* 0.809 0.368 0.57 0.291 0.0271 0.0372 0.132 0.239 0.0117 0.0909
@@
Table 2: Quantitative evaluation of THUNDER, other models and ablations. (1) Mesh-to-speech input space. All of the mesh-to-speech variants are applicable. THUNDER with mouth2s results in the best lip-sync performance (LVE, L-CCC, L-PCC), which is why we select it as the final THUNDER model. exp2s is applicable, too as its scores best on DTW and has good diversity (S-DIV-U), while face2s scores the worst on diversity (S-DIV-U, T-DIV-U). (2) Trainable audio encoder. THUNDER-T results in improved lip-sync metrics (LVE, L-CCC, L-PCC and DTW), but at the slight expense of diversity of outputs (mainly upper face diversity S-DIV-U and T-DIV-U) compared to THUNDER which does not train the audio encoder. (3) Mesh-To-Speech applied to other methods. The application of mesh-to-speech to our reimplementations of Media2Face* (diffusion with image conditioning), DiffPoseTalk* (diffusion with style feature conditioning) and FlameFormer (deterministic with subject ID conditioning) improves lip-sync metrics at slight expense of expression diversity. Please note that DiffPoseTalk* and Media2Face* (both with and without mesh-to-speech) have a considerably lower diversity than THUNDER. This happens due to the additional conditioning inputs which narrow down the output distribution space. This also manifests itself with better lip-sync performance.
@@
Ablation study. We ablate our design decisions. (1) The input space for the mesh-to-speech loss, (2) the trainability of the audio encoder (Tab. 2) , and (3) the weight of the mesh-to-speech loss (see Sec. 5 in Sup. Mat. and video).
Mesh-to-speech loss with other methods. Tab. 2 shows that mesh-to-speech is applicable to other methods. We re-implement three methods that incorporate additional conditioning mechanisms and train them with and without our mesh-to-speech loss. Our new loss improves performance on the lip-sync metrics. For visual results, refer to Sup. Video. Comparison to additional methods. Tab. 7 in Sup. Mat. we also include a comparison of THUNDER and other (original) implementations recent methods.
@@
Disentanglement effect. To test whether mesh-to-speech helps preserve the effect of good lip-sync in the presence of other editing conditions, we additionally run our reimplementations of Media2Face with conditioning inputs extracted from emotional images. See Sec. 7.2.1 in Sup. Mat.
Other datasets. To verify, whether THUNDER is dataset-agnostic, we also trained and tested the method on TFHP. We report the results in Sec. 7.2.4 in Sup. Mat..
@@
Perceptual study. We conduct a perceptual study in which we evaluate THUNDER and the effect of mesh-to-speech. Specifically, we compare to the deterministic FlameFormer, THUNDER with and without mesh-to-speech, and p-GT. The participants are shown two videos side-by-side and asked to rate animations on a five-point Likert scale (strong/weak preference for A or B or indifferent). We ask the participants to rate three aspects of the animation: lip-sync quality, dynamism, and realism. Fig. 4 reports the results for THUNDER and Fig. 5 for THUNDER-T. For the remaining results of the study and further discussion, see Sec. 13 in Sup. Mat.
Figure 4: Perceptual study of THUNDER. We compare THUNDER-F with methods having both a trainable audio encoder (THUNDER-T w/o m2s) and frozen encoders (FlameFormer-F, THUNDER-F w/o m2s), and GT. The participants prefer THUNDER-F’s lip-sync over that of other models, but does not quite reach the fidelity of GT. Remarkably, the participants also have slight preference for THUNDER-F over GT, which suggests that the application of mesh-to-speech helps THUNDER saturate the quality of GT.
@@
Figure 5: Perceptual study of THUNDER-T. We compare THUNDER-T with methods having both a trainable audio encoders (FlameFormer-T, THUNDER-T w/o m2s) and GT. Models with trainable backbones, achieve good lip-sync than their frozen counterparts, but THUNDER-T is still preferred over the other models.
@@ 5.4 Limitations
THUNDER has the following limitations. The absence of teeth and tongue inherently creates ambiguity, making it more difficult to produce high quality audio compared to silent video-to-speech. Further, pseudo-GT is not a perfect reconstruction but has occasional artifacts or inaccuracies. Also, our current mesh-to-speech architecture is fully deterministic and hence is not completely capable of capturing the prediction ambiguities. All of these may result in suboptimal effects of the audio cycle-consistency loss. However, regardless of these shortcomings, our experiments demonstrate the benefits of our method. Employing a large-scale dataset of high quality 3D scans, a face model that models teeth and tongue, and upgrade mesh-to-speech to a stochastic model are likely to boost the benefits of the mesh-to-speech loss. Finally, the inference process of the diffusion model is currently computationally intensive. Future work should address this by employing more efficient solvers such as DPM++ [lu2022dpmpp].
@@ 6 Conclusion
[…]
Supplementary Material
7 Additional Experiments
7.1 Mesh-to-Speech
Per-Dataset Performance Analysis. The performance of both mesh-to-speech and video-to-speech models varies dramatically depending on the dataset. Datasets with limited number of utterances and vocabulary such as RAVDESS or GRID are less challenging, and the models can regress relatively accurately, sometimes to the point of perfect intelligibility. On the more challenging TCD-TIMIT dataset, which contains more complex vocabulary, words are often missed. However, despite the relatively high word error rate (WER) on TCD-TIMIT, the produced sounds, remain very realistic, given the animation. Please see Tab. 3 for per-dataset breakdown and refer to the supplementary video for an audible comparison.
@@ 7.1 Mesh-to-Speech
7.2 Talking Heads
7.2.1 Disentanglement